OpenStudy (anonymous):

3/y^2-36 - 2/y^2+4y-12

jhonyy9 (jhonyy9):

3/((y-6)(y+6)) - 2/((y+6)(y-2)) = (3(y-2)-2(y-6))/((y-6)(y+6)(y-2)) = = (3y-6-2y+12)/((y-6)(y+6)(y-2))=(y+6)/((y-6)(y+6)(y-2)) = 1/((y-6)(y-2))

OpenStudy (lalaly):

\[(\frac{3}{(y-6)(y+6)} \times \frac{y-2}{y-2}) - (\frac{2}{(y+6)(y-2)} \times \frac{y-6}{y-6})\]\[=\frac{3(y-2)-2(y-6)}{(y-6)(y-2)(y+6)} = \frac{3y-6-2y+12}{(y-6)(y-2)(y+6)}\]\[=\frac{\cancel{y+6}}{(y-6) \cancel{(y+6)}(y-2)}= \frac{1}{(y-6)(y-2)}\]

jhonyy9 (jhonyy9):

i have wrote exactly

jhonyy9 (jhonyy9):

thank you lalay

OpenStudy (lalaly):

:):):)